OpenStudy (anonymous):

3x - (1/2) = 2x + (7/4)

OpenStudy (anonymous):

Well then: 3x - 2x = 7/4 + 1/2 x = 7/4 + 2/4 x = 9/4.

OpenStudy (anonymous):

okay thanks and thanks for showing the steps